# Physics-informed Kolmogorov–Arnold Network with Chebyshev Polynomials for Fluid Mechanics

### v2

@@ Abstract
Solving partial differential equations (PDEs) is essential in scientific forecasting and fluid dynamics. Traditional approaches often incur expensive computational costs and trade-offs in efficiency and accuracy. Recent deep neural networks improve the accuracy but require high-quality training data. Physics-informed neural networks (PINNs) effectively integrate physical laws to reduce the data reliance in limited sample scenarios. A novel machine-learning framework, Chebyshev physics-informed Kolmogorov–Arnold network (ChebPIKAN), is proposed to integrate the robust architectures of Kolmogorov–Arnold networks (KAN) with physical constraints to enhance the calculation accuracy of PDEs for fluid mechanics. We study the fundamentals of KAN, take advantage of the orthogonality of Chebyshev polynomial basis functions in spline fitting, and integrate physics-informed loss functions that are tailored to specific PDEs in fluid dynamics, including Allen–Cahn equation, nonlinear Burgers equation, Helmholtz equations, Kovasznay flow, cylinder wake flow, and cavity flow. Extensive experiments demonstrate that the proposed ChebPIKAN model significantly outperforms the standard KAN architecture in solving various PDEs by effectively embedding essential physical information. These results indicate that augmenting KAN with physical constraints can alleviate the overfitting issues of KAN and improve the extrapolation performance. Consequently, this study highlights the potential of ChebPIKAN as a powerful tool in computational fluid dynamics and propose a path toward fast and reliable predictions in fluid mechanics and beyond.
@@
In recent years, machine learning (ML) techniques have quickly gained attention and been rapidly developed in fluid mechanics research due to their minimal requirements for physical information and fast response capabilities. At present, in mathematics and practical engineering, deep neural networks have been applied to solving partial differential equations due to their excellent nonlinear fitting and interpolation capabilities. Srinivasan et al.Srinivasan et al. (2019) evaluated the ability of deep neural networks to predict temporal turbulent flows, generated training data using a nine-equation shear flow model, and tested the multi-layer perceptron (MLP) and long short-term memory (LSTM) networks. The results indicate that LSTM is better than MLP in predicting turbulence statistics and dynamic behaviors since LSTM can use the sequential characters of data with relatively small relative errors. This exploratory study established the foundation for the future development of accurate and effective data-driven sub-grid scale models to apply them to more complex large-eddy simulations of wall turbulence. Bhatnagar et al.Bhatnagar et al. (2019) proposed a convolutional-neural-network (CNN)-based model to predict the airfoil flow field, where Reynolds-averaged Navier Stokes (RANS) solutions on airfoil profiles were used as training data. This model automatically extracts mapping features with minimal human supervision and can more rapidly predict the velocity and pressure fields than RANS solvers while enhancing prediction capabilities through specific convolution operations and parameter sharing. Santos et al.Santos et al. (2020) introduced PoreFlow Net, which is a three-dimensional (3D) convolutional neural network architecture that significantly accelerates the flow field prediction process by extracting the spatial relationship between porous media morphology and fluid velocity field. This approach alleviates the need for costly numerical simulations and demonstrates the successful application of physics-based machine learning models in digital rock physics. Kim et al.Kim and Lee (2020) used a CNN framework to predict the local heat flux through direct-numerical-simulation data and demonstrated that the accuracy was preserved even at higher Reynolds numbers. These results enhanced the understanding of turbulence heat transfer and provided a novel method for turbulence modeling. W. Chen et al.Chen et al. (2021) proposed a physics-informed machine learning framework for efficient reduced-order modeling of parametric PDEs. Through performance comparisons with traditional methods, the framework demonstrates significant potential for the efficient reduced-order solution of PDEs. Erichson et al.Erichson et al. (2020) proposed a data-driven method that used shallow neural networks to reconstruct flow fields from limited measurement data, which relieved the requirement for complex preprocessing. Kashefi et al.Kashefi, Rempe, and Guibas (2021) proposed a deep learning framework based on the PointNet architecture to predict the flow field in irregular domains, where point cloud inputs were used to embed the spatial positions into CFD quantities. This method effectively preserves geometric information and trains much faster than traditional CFD solvers without decreasing the prediction accuracy.
@@
Based on the excellent capability of PINN in solving physical equations, numerous improved network architectures based on PINN have been successively proposed. Inspired by the application of neural networks in computer vision, Levi D et al.McClenny and Braga-Neto (2023) designed self-adaptive weights for PINN, enabling it to autonomously identify and focus on difficult fitting regions during the training process. Through comparative validation, their approach demonstrated superior performance over traditional PINN. S. Lin et al.Lin and Chen (2022) appended a completely new neural network after the original PINN, incorporating measurements of conservative quantities into the mean squared error loss. This approach demonstrated higher accuracy and generalization ability in standard fluid dynamics cases. Yuan L et al.Yuan et al. (2022) proposed auxiliary physics-informed neural networks (A-PINNs), where auxiliary output variables are introduced to represent the numerical integrals in the governing equations and can be automatically differentiated instead of integral operators to avoid the limitation of using integral discretization. Using this neural network to solve nonlinear integral differential equations yields better accuracy than PINN. Inspired by the finite element method, Rezaei S et al.Rezaei et al. (2022) sought to improve the performance of the existing PINNs by advocating the use of spatial gradients of the principal variables as the output of the separable neural network. By properly designing the network architecture in PINN, deep learning models can solve unknown problems in heterogeneous domains without initial data from other sources. Recent work by S. Zhao et al.Zhao et al. (2025) directly encodes the Large-Eddy Simulation (LES) equations into neural operators for three-dimensional incompressible turbulent flow simulation, with the resulting LESnets model demonstrating accuracy comparable to conventional methods while exhibiting strong generalization capabilities to unseen flow states. The physics-informed neural-network-based topology optimization (PINNTO), combines PINN with topology optimization, and solves several inherent defects due to the addition of PINN Jeong et al. (2023). The accuracy of PINN also highly depends on the selection and distribution of training sample points. Y Liu et al.Liu et al. (2024a) proposed a novel adaptive sampling algorithm called expected-improvement residual-based adaptive refinement (EI-RAR), which combines an attention mechanism with sample point generation using a new expected-improvement function. This approach addresses the limitation of traditional adaptive sampling algorithms, which typically focus only on sample points in the solution domain. Furthermore, the expected-improvement gradient-adaptive sampling algorithm, which integrates residual gradient information, was proposed to enhance the nonlinear fitting accuracy of PINN in discontinuous solution regions. J. Bai et al.Bai et al. (2023) extended the PINN algorithm to the solid mechanics and proposed the LSWR loss function based on the least square weighted residual (LSWR) method. Comparisons with energy-based and collocation loss functions have demonstrated the potential of the LSWR loss function in accurately predicting stress and displacement fields. H. Gao et al.Gao, Sun, and Wang (2021) employed elliptic coordinate mapping to achieve coordinate transformation between irregular physical domains and regular reference domains, to leverage powerful classic CNN backbones for solving steady-state PDEs. The designed physics-informed geometry-adaptive convolutional neural networks (GhyGeoNet) has been demonstrated to have significant research value. L Yang et al.Liu et al. (2023) research on PINN transfer learning tasks indicates that in the future, PINN would be more easily applied to new physical cases, and the development of physical fields would to some extent, depend on its progress.
@@
Although numerous neural network architectures have been developed to enhance the performance of PINN for different problems, their fundamental structure remains rooted in the original PINN framework. PINN encounters the issues of gradient vanishing during the training process and is highly sensitive to the selection of various hyperparameters. However, extensive parameters require manual tuning, and there is currently no established optimal parameter selection guideline for PINN Faroughi et al. (2024). Inspired by Kolmogorov–Arnold representation theorem, Kolmogorov–Arnold networks (KANs) have fundamentally transformed the traditional MLP architecture by fitting the weight coefficients of the learning activation function with learnable univariate function spline curves at the edges instead of using fixed activation functions at the nodes as in conventional multilayer perceptron (MLP) Liu et al. (2024b). Thus, KANs do not require the use of any nonlinear activation functions, are closer to physical laws, and better align with human intuition. In traditional neural network development, the nonlinear mapping between inputs and outputs is often an obscure mapping that lacks interpretability, and the use of nonlinear activation functions cannot be easily explained from a physical perspective. Thus, traditional neural networks function as "black boxes" and make their results inherently difficult to interpret and less convincing de Francca (2018). Since the learning process of the spline fitting of KANs is grounded in mathematical theorems, the network architecture of KANs inherently possesses physical significance. From a machine learning perspective, incorporating physical information as constraints in neural networks is a significant development trend. This approach offers several advantages such as introducing physical constraints to the loss function to enhance the physical interpretability of neural networks Willard et al. (2022). However, KANs may still require a substantial amount of training data to achieve ideal results since there are no physical information constraints. Inspired by the concepts of PINN, incorporating machine learning techniques with prior knowledge such as physical information is a critical research approach Li, Du, and Martins (2022). Although KAN has inherent physical significance, we believe that it is essential to incorporate physical information constraints into KAN, which is similar to the enhancements in PINN compared with conventional deep neural networks, and a similar approach has also been employed by L Yang et al.Yang, Meng, and Karniadakis (2021) in their study of partial differential equations with noisy data, where they combine physical information with Bayesian Neural Networks (BNN) to achieve improved predictive accuracy. In this study, we propose a novel Kolmogorov–Arnold network architecture based on Chebyshev polynomials, i.e., Chebyshev physics-informed Kolmogorov–Arnold network (ChebPIKAN). The proposed ChebPIKAN framework adopts the Kolmogorov–Arnold network as the baseline network architecture and incorporates essential physical information based on PINN. KAN can adaptively learn the nonlinear activation functions in comparison with the classical neural network and automatically prune the basis functions and weight coefficients to determine the optimal network architecture. The proposal of KAN fundamentally solves the problem of poor generalization caused by the global optimization of traditional neural networks, since the spline curves learned by KAN are essentially local, and the fitted splines learned by the network are independent Liu et al. (2024b). In recent research findings, KAN has also been demonstrated its effectiveness in the field of physics Koenig, Kim, and Deng (2024); Wang et al. (2024). The proposed ChebPIKAN framework uses KAN as its foundational architecture, inherits the advantages of the KAN structure, and achieves enhanced performance due to the incorporation of physical information.
@@ II.1 Kolmogorov–Arnold Networks
where $\phi_{q,p}:[0,1]\rightarrow\mathbf{R}$ and $\Phi_{q}:\mathbf{R}\rightarrow\mathbf{R}$. The aim is to identify suitable univariate functions $\phi_{q,p}$ and $\Phi$. $\phi(x)$ represents the operations of the KAN layer’s node, and the input of the KAN layer is $\bm{x}$, which leads to the mathematical formula for the $l-th$ KAN layer
@@
where $\Phi_{l}$ is the operation matrix for the $l-th$ KAN layer, which contains learnable parameters.
@@ II.2 Chebyshev Polynomial Basis Functions
where $\omega$ is the learnable weight coefficient of neural network. $x$ is the input value of function $\phi$ instead of a vector that contains multiple arguments, as shown in Section 2.1. $spline(x)$ is modeled as a linear combination of B-$splines$ functions, and the activation function is defined by
@@
The goal of neural networks is to learn appropriate weight coefficients $\omega$ to approximate the target function. However, the B-$spline$ function is not an orthogonal basis function and has a linear correlation, which may cause multi-collinearity issues during the fitting process. Multi-collinearity can make the parameter estimation unstable, increase the variance of the fitting results, and make it difficult to explain the contributions of individual basis function. In practical applications, high-frequency oscillations (such as the Runge phenomenon) likely occur. To address this issue, we use Chebyshev polynomials, which provide orthogonality and enhance the interpolation performance.
Chebyshev polynomials consist of two polynomial sequences related to cosine and sine functions, denoted by $T_{n}(x)$ and $U_{n}(x)$. They can be defined in several equivalent manners. The $T_{n}(x)$ Chebyshev polynomial is derived by the recursive formula
@@
with $x\in[-1,1]$ and $T_{n}(x)\in[-1,1]$. Since the initial input $x$ or the output after a single KAN layer can easily exceed the value range of the Chebyshev polynomial, we apply the tanh activation function to the output of the KAN layer and the first input $x_{0}$ to limit it within the range of $[-1,1]$. Therefore, the input of each KAN layer undergoes additional $\tanh$ activation
@@
where $n\geq 1$ and $cheb(\cdot)$ are the Chebyshev polynomial in Eq. (5).

### v3

@@ Abstract
Solving partial differential equations (PDEs) is essential in scientific forecasting and fluid dynamics. Traditional approaches often incur expensive computational costs and trade-offs in efficiency and accuracy. Recent deep neural networks have improved the accuracy but require high-quality training data. Physics-informed neural networks (PINNs) effectively integrate physical laws to reduce the data reliance in limited sample scenarios. A novel machine-learning framework, Chebyshev physics-informed Kolmogorov–Arnold network (ChebPIKAN), is proposed to integrate the robust architectures of Kolmogorov–Arnold networks (KAN) with physical constraints to enhance the calculation accuracy of PDEs for fluid mechanics. We study the fundamentals of KAN, take advantage of the orthogonality of Chebyshev polynomial basis functions in spline fitting, and integrate physics-informed loss functions that are tailored to specific PDEs in fluid dynamics, including Allen–Cahn equation, nonlinear Burgers equation, Helmholtz equations, Kovasznay flow, cylinder wake flow, and cavity flow. Extensive experiments demonstrate that the proposed ChebPIKAN model significantly outperforms the standard KAN architecture in solving various PDEs by effectively embedding essential physical information. These results indicate that augmenting KAN with physical constraints can alleviate the overfitting issues of KAN and improve the extrapolation performance. Consequently, this study highlights the potential of ChebPIKAN as a powerful tool in computational fluid dynamics and propose a path toward fast and reliable predictions in fluid mechanics and beyond.
@@
In recent years, machine learning (ML) techniques have quickly gained attention and have been rapidly developed in fluid mechanics research due to their minimal requirements for physical information and fast response capabilities. At present, in mathematics and practical engineering, deep neural networks have been applied to solving partial differential equations due to their excellent nonlinear fitting and interpolation capabilities. Srinivasan et al.Srinivasan et al. (2019) evaluated the ability of deep neural networks to predict temporal turbulent flows, generated training data using a nine-equation shear flow model, and tested the multi-layer perceptron (MLP) and long short-term memory (LSTM) networks. The results indicate that LSTM is better than MLP in predicting turbulence statistics and dynamic behaviors since LSTM can use the sequential characters of data with relatively small relative errors. This exploratory study established the foundation for the future development of accurate and effective data-driven sub-grid scale models to apply them to more complex large-eddy simulations of wall turbulence. Bhatnagar et al.Bhatnagar et al. (2019) proposed a convolutional-neural-network (CNN)-based model to predict the airfoil flow field, where Reynolds-averaged Navier Stokes (RANS) solutions on airfoil profiles were used as training data. This model automatically extracts mapping features with minimal human supervision and can more rapidly predict the velocity and pressure fields than RANS solvers while enhancing prediction capabilities through specific convolution operations and parameter sharing. X Zhang et al.Zhang et al. (2024) proposed a Residual Graph Convolutional Network for Initial Flow Field Setting (RGCN-IFS) that provides high-quality initial flow fields for CFD simulations, effectively bridging the accuracy gap between intelligent surrogate models and conventional CFD approaches while maintaining computational efficiency. Santos et al.Santos et al. (2020) introduced PoreFlow Net, which is a three-dimensional (3D) convolutional neural network architecture that significantly accelerates the flow field prediction process by extracting the spatial relationship between porous media morphology and fluid velocity field. This approach alleviates the need for costly numerical simulations and demonstrates the successful application of physics-based machine learning models in digital rock physics. Kim et al.Kim and Lee (2020) used a CNN framework to predict the local heat flux through direct-numerical-simulation data and demonstrated that the accuracy was preserved even at higher Reynolds numbers. These results enhanced the understanding of turbulence heat transfer and provided a novel method for turbulence modeling. W. Chen et al.Chen et al. (2021) proposed a physics-informed machine learning framework for efficient reduced-order modeling of parametric PDEs. Through performance comparisons with traditional methods, the framework demonstrates significant potential for the efficient reduced-order solution of PDEs. Erichson et al.Erichson et al. (2020) proposed a data-driven method that used shallow neural networks to reconstruct flow fields from limited measurement data, which relieved the requirement for complex preprocessing. Kashefi et al.Kashefi, Rempe, and Guibas (2021) proposed a deep learning framework based on the PointNet architecture to predict the flow field in irregular domains, where point cloud inputs were used to embed the spatial positions into CFD quantities. This method effectively preserves geometric information and trains much faster than traditional CFD solvers without decreasing the prediction accuracy.
@@
Based on the excellent capability of PINN in solving physical equations, numerous improved network architectures based on PINN have been successively proposed. Inspired by the application of neural networks in computer vision, Levi D et al.McClenny and Braga-Neto (2023) designed self-adaptive weights for PINN, enabling it to autonomously identify and focus on difficult fitting regions during the training process. Through comparative validation, their approach demonstrated superior performance over traditional PINN. S. Lin et al.Lin and Chen (2022) appended a completely new neural network after the original PINN, incorporating measurements of conservative quantities into the mean squared error loss. This approach demonstrated higher accuracy and generalization ability in standard fluid dynamics cases. Yuan L et al.Yuan et al. (2022) proposed auxiliary physics-informed neural networks (A-PINNs), where auxiliary output variables are introduced to represent the numerical integrals in the governing equations and can be automatically differentiated instead of integral operators to avoid the limitation of using integral discretization. Using this neural network to solve nonlinear integral differential equations yields better accuracy than PINN. Inspired by the finite element method, Rezaei S et al.Rezaei et al. (2022) sought to improve the performance of the existing PINNs by advocating the use of spatial gradients of the principal variables as the output of the separable neural network. By properly designing the network architecture in PINN, deep learning models can solve unknown problems in heterogeneous domains without initial data from other sources. Recent work by S. Zhao et al.Zhao et al. (2025) directly encodes the Large-Eddy Simulation (LES) equations into neural operators for three-dimensional incompressible turbulent flow simulation, with the resulting LESnets model demonstrating accuracy comparable to conventional methods while exhibiting strong generalization capabilities to unseen flow states. The physics-informed neural-network-based topology optimization (PINNTO), combines PINN with topology optimization, and solves several inherent defects due to the addition of PINN Jeong et al. (2023). The accuracy of PINN also highly depends on the selection and distribution of training sample points. Y Liu et al.Liu et al. (2024a) proposed a novel adaptive sampling algorithm called expected-improvement residual-based adaptive refinement (EI-RAR), which combines an attention mechanism with sample point generation using a new expected-improvement function. This approach addresses the limitation of traditional adaptive sampling algorithms, which typically focus only on sample points in the solution domain. Furthermore, the expected-improvement gradient-adaptive sampling algorithm, which integrates residual gradient information, was proposed to enhance the nonlinear fitting accuracy of PINN in discontinuous solution regions. W Zhou et al.Zhou, Miwa, and Okamoto (2024) proposed an enhanced PINN framework that synergistically integrates residual-based adaptive sampling, adaptive loss weighting, and differential evolution optimization algorithms, demonstrating its effectiveness through several benchmark test cases. J. Bai et al.Bai et al. (2023) extended the PINN algorithm to the solid mechanics and proposed the LSWR loss function based on the least square weighted residual (LSWR) method. Comparisons with energy-based and collocation loss functions have demonstrated the potential of the LSWR loss function in accurately predicting stress and displacement fields. H. Gao et al.Gao, Sun, and Wang (2021) employed elliptic coordinate mapping to achieve coordinate transformation between irregular physical domains and regular reference domains, to leverage powerful classic CNN backbones for solving steady-state PDEs. The designed physics-informed geometry-adaptive convolutional neural networks (GhyGeoNet) has been demonstrated to have significant research value. Recently, S Wang et al.Wang et al. (2025a) successfully employed PINNs to simulate three-dimensional fully turbulent flows at high Reynolds numbers, marking a significant step toward more flexible and continuous computational fluid dynamics. L Yang et al.Liu et al. (2023) research on PINN transfer learning tasks indicates that in the future, PINN would be more easily applied to new physical cases, and the development of physical fields would to some extent, depend on its progress.
@@
Although numerous neural network architectures have been developed to enhance the performance of PINN for different problems, their fundamental structure remains rooted in the original PINN framework. PINN encounters the issues of gradient vanishing during the training process and is highly sensitive to the selection of various hyperparameters. However, extensive parameters require manual tuning, and there is currently no established optimal parameter selection guideline for PINN Faroughi et al. (2024). Inspired by Kolmogorov–Arnold representation theorem, Kolmogorov–Arnold networks (KANs) have fundamentally transformed the traditional MLP architecture by fitting the weight coefficients of the learning activation function with learnable univariate function spline curves at the edges instead of using fixed activation functions at the nodes as in conventional multilayer perceptron (MLP) Liu et al. (2024b). Thus, KANs require no additional nonlinear activation functions between network nodes. This design renders KANs more physically consistent and intuitively aligned with human cognition. In traditional neural network development, the nonlinear mapping between inputs and outputs is often an obscure mapping that lacks interpretability, and the use of nonlinear activation functions cannot be easily explained from a physical perspective. Thus, traditional neural networks function as "black boxes" and make their results inherently difficult to interpret and less convincing de Francca (2018). Since the learning process of the spline fitting of KANs is grounded in mathematical theorems, the network architecture of KANs inherently possesses physical significance. From a machine learning perspective, incorporating physical information as constraints in neural networks is a significant development trend. This approach offers several advantages such as introducing physical constraints to the loss function to enhance the physical interpretability of neural networks Willard et al. (2022). However, KANs may still require a substantial amount of training data to achieve ideal results since there are no physical information constraints. Inspired by the concepts of PINN, incorporating machine learning techniques with prior knowledge such as physical information is a critical research approach Li, Du, and Martins (2022). Although KAN has inherent physical significance, we believe that it is essential to incorporate physical information constraints into KAN, which is similar to the enhancements in PINN compared with conventional deep neural networks, and a similar approach has also been employed by L Yang et al.Yang, Meng, and Karniadakis (2021) in their study of partial differential equations with noisy data, where they combine physical information with Bayesian Neural Networks (BNN) to achieve improved predictive accuracy. In this study, we propose a novel Kolmogorov–Arnold network architecture based on Chebyshev polynomials, i.e., Chebyshev physics-informed Kolmogorov–Arnold network (ChebPIKAN). The proposed ChebPIKAN framework adopts the Kolmogorov–Arnold network as the baseline network architecture and incorporates essential physical information based on PINN. KAN can adaptively learn the nonlinear activation functions in comparison with the classical neural network and automatically prune the basis functions and weight coefficients to determine the optimal network architecture. The proposal of KAN fundamentally solves the problem of poor generalization caused by the global optimization of traditional neural networks, since the spline curves learned by KAN are essentially local, and the fitted splines learned by the network are independent Liu et al. (2024b). In recent research findings, KAN has also been demonstrated its effectiveness in the field of physics Koenig, Kim, and Deng (2024); Wang et al. (2025b). Specifically, the Kolmogorov–Arnold-Informed Neural Network (KINN), proposed by Y Wang et al.Wang et al. (2025b), provides a novel approach for solving both forward and inverse problems of PDEs. Tests on multiple PDE benchmarks demonstrate that KINN surpasses traditional MLP methods in accuracy and convergence speed, offering a more efficient and accurate machine learning framework for PDE solutions. The proposed ChebPIKAN framework uses KAN as its foundational architecture, inherits the advantages of the KAN structure, and achieves enhanced performance due to the incorporation of physical information.
@@ II.1 Kolmogorov–Arnold Networks
where $\phi_{q,p}:[0,1]\rightarrow\mathbf{R}$ and $\Phi_{q}:\mathbf{R}\rightarrow\mathbf{R}$. The aim is to identify suitable univariate functions $\phi_{q,p}$ and $\Phi$. $\phi(x)$ represents the operations of the KAN layer’s node, and the input of the KAN layer is $\bm{x}$, which leads to the mathematical formula for the lthl-th KAN layer
@@
where $\Phi_{l}$ is the operation matrix for the lthl-th KAN layer, which contains learnable parameters.
@@ II.2 Chebyshev Polynomial Basis Functions
where $\omega$ is the learnable weight coefficient of neural network. $x$ is the input value of function $\phi$ instead of a vector that contains multiple arguments, as shown in Section II.A, $spline(x)$ is modeled as a linear combination of B-splinessplines functions, and the activation function is defined by
@@
The goal of neural networks is to learn appropriate weight coefficients $\omega$ to approximate the target function. However, the B-splinespline function is not an orthogonal basis function and has a linear correlation, which may cause multi-collinearity issues during the fitting process. Multi-collinearity can make the parameter estimation unstable, increase the variance of the fitting results, and make it difficult to explain the contributions of individual basis function. In practical applications, high-frequency oscillations (such as the Runge phenomenon) likely occur. To address this issue, we use Chebyshev polynomials, which provide orthogonality and enhance the interpolation performance.
Chebyshev polynomials consist of two polynomial sequences related to cosine and sine functions, denoted by $T_{n}(x)$ and $U_{n}(x)$. They can be defined in several equivalent manners. The $T_{n}(x)$ Chebyshev polynomials are derived by the recursive formula
@@
with $x\in[-1,1]$ and $T_{n}(x)\in[-1,1]$. Since the initial input $x$ or the output after a single KAN layer can easily exceed the value range of the Chebyshev polynomials, we apply the tanh normalization function to the output of the KAN layer and the first input $x_{0}$ to limit it within the range of [1,1][-1,1]. Therefore, the input of each KAN layer undergoes additional $\tanh$ normalization
@@
where $n\geq 1$ and $cheb(\cdot)$ are the Chebyshev polynomials in Eq. (5).